A boy can row a boat at a constant rate of 3mi/hr in still water. He rows up-stream for 20 minutes and then rows down-stream, returning to his starting point in another 15 minutes. Find the total distance traveled.
Answer by stanbon(75887)

A boy can row a boat at a constant rate of 3mi/hr in still water. He rows up-stream for 20 minutes and then rows down-stream, returning to his starting point in another 15 minutes. Find the total distance traveled.
------------------------------------
Upstream DATA:
time = 1/5 hr ; rate = 3+c ; distance = rt = (1/5)(3+c) miles
------------------------------------
Downstream DATA:
time = 1/4 hr ; rate = 3-c ; distance = rt = (1/4)(3-c) miles
------------------------------------
Equation:
Up distance = Down distance
(1/5)(3+c) = (1/4)(3-c)
Multiply both sides by 20 to get:
4(3+c) = 5(3-c)
12+4c = 15-5c
9c = 3
c = 1/3 mph (speed of the current
=========================================
Cheers,
Stan H.